OpenStudy (anonymous):

Simplify This Expression. (7x^5)(8x) =

OpenStudy (anonymous):

7x^5 * 8x = 56 * x^5 * x =

OpenStudy (anonymous):

now the last step looks like the question that ive answered with x instead of 2

OpenStudy (anonymous):

x is a common base so..

OpenStudy (anonymous):

so..

OpenStudy (anonymous):

a^b * a^c = a^(b+c)

OpenStudy (anonymous):

so you have here: 56 * x^5 * x = 56 * x^5 * x^1 = ..

OpenStudy (anonymous):

is there an easier way.

OpenStudy (anonymous):

i want you to write the final answer

OpenStudy (anonymous):

56 * x^5 * x^1 = 56 * x^(5+1) = 56x^6